OpenStudy (anonymous):

Which ordered pair is NOT a solution of the inequality, 3x - 2y is greater than or equal to. 12 A. (-3, -10.5) B. (-1, -7) C. (0, -6) D. (2, -3)

OpenStudy (anonymous):

@shadeyghost

OpenStudy (shadeyghost):

im thinking

OpenStudy (anonymous):

@pooja195

OpenStudy (shadeyghost):

wow, i can't think

OpenStudy (shadeyghost):

i need a graph or something

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

look up graph

OpenStudy (anonymous):

Just plug in the points.

OpenStudy (anonymous):

\(3x - 2y \ge 12\) \((-3, -10.5)\) \(3(-3) - 2(-10.5) \ge 12\) Simplify and see if it's true or not.

OpenStudy (anonymous):

Can you do that? @kiker

OpenStudy (anonymous):

ya hold on im doing an equations

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Actually I think that's true.

OpenStudy (anonymous):

I get: \(12 \ge 12\) Greater than OR equal to

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

You put in 10.5..it's supposed to be -10.5.

OpenStudy (anonymous):

But it's asking which is NOT a solution..

OpenStudy (anonymous):

so not a

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

No, it's not A, because A is a solution..and it's looking for one that isn't a solution.

OpenStudy (anonymous):

Okay, simplify this one now: \(3(-1) - 2(-7) \ge 12\)

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Yep, so that's your answer.

OpenStudy (anonymous):

thanks again